OpenStudy (anonymous):

If 2a+b=7 and b+2c=23, what is the average of a, b, and c?

OpenStudy (anonymous):

Adding the 2 equations, i get- \[2a+2b+2c=7+23\\2(a+b+c)=30\\a+b+c=15.\text{Average of a, b, c = }\frac{a+b+c}{2}\;=\;\frac{15}{2}\;=\;7.5\]